A 1.35-kg block of wood sits at the edge of a table, 0.730 m above the floor. A 1.15×10−2-kg bullet moving horizontally with a speed of 745 m/s embeds itself within the block.What horizontal distance does the block cover before hitting the ground?
Answer #1

Gravitational acceleration = g = -9.81 m/s2

Mass of the bullet = m1 = 1.15 x 10-2 kg

Mass of the block of wood = m2 = 1.35 kg

Initial velocity of the bullet = V1 = 745 m/s

Initial velocity of the block of wood = V2 = 0 m/s

Velocity of the block of wood and bullet after the collision = V3

By conservation of linear momentum,

m1V1 + m2V2 = (m1 + m2)V3

(1.15x10-2)(745) + (1.35)(0) = (1.15x10-2 + 1.35)V3

V3 = 6.293 m/s

The block will leave the table horizontally therefore,

Horizontal velocity of the block when it leaves the table = V3x = 6.293 m/s

Vertical velocity of the block when it leaves the table = V3y = 0 m/s

Height of the table = H = 0.73 m

Time taken by the block to hit the floor = T

The displacement of the block in the vertical direction when it hits the floor is downwards therefore it is negative.

-H = V3yT + gT2/2

-0.73 = (0)T + (-9.81)T2/2

T = 0.386 sec

Horizontal distance covered by the block = R

There is no horizontal force acting on the block therefore the horizontal velocity of the block remains constant.

R = V3xT

R = (6.293)(0.386)

R = 2.43 m

Horizontal distance the block covers before hitting the ground = 2.43 m